Calculate the present value of each cashflow using a discount rate of 7%. Which do you most prefer most?

Show and explain all supporting calculations!

Cashflow A: receive $60 today and then receive $60 in four years.

Cashflow B: receive $12 every year, forever, starting today.

Cashflow C: pay $50 every year for five years, with the first payment being next year, and then subsequently receive $30 every year for 20 years.

Cashflow D: receive $9 every other year, forever, with the first payment being next year.
Answer #1

Ans-A- PV= 60+60/ (1+0.07)^4

=60+60/1.310796

=60+45.77371

=$105.77

B- PV of perpetuity= 12+12/0.07

=12+171.43

=$183.43

C-PV= -50* PV factor of annuity @7%, n= 5+(30*PV factor of annuity @7% for 20 years)* PV factor @ 7% for 5 years

=-50/0.07*(1-1.07^-5)+ 30/0.07*(1-1.07^-20)*1.07^-5

=$21.59

D-PV of perpetuity= 9/(1.07^2-1)*1.07

=$66.46

The PV of cashflow B is largest and should be most preferred.